Mercy invested part of her money at 8% per year and twice amount at 5% per year. If the return last year in the two investment was 4200 how much did mercy invest at each rate
Answer by Theo(13342)

let x = the amount of money she invested at 8%.
then 2 * x = the amount of money she invested at 5%.

.08 * x + .05 * 2 * x = 4200.

simplify to get .08 * x + .10 * x = 4200.

combine like terms to get .18 * x = 4200.

divide both sides of this equation by .18 to get x = 23,333.33 rounded to 2 decimal places.

2x is therefore equal to 46,666.67 rounded to 2 decimal places.

you get .08 * 23,333.33 + .05 * 46,666.67 = 4200.

i used the unrounded numbers to do the calculations.

your solution is that she invested 23,333.33 at 8% and 46,666.67 at 5%.